OpenStudy (anonymous):

three times the smallest of 3 consecutive even integers plus the sum of the other 2 is 66. find the numbers. Please help, thanks

OpenStudy (anonymous):

12 14 16

OpenStudy (anonymous):

assume numbers are x x+2 x+4 3x+(x+2)+(x+4)=66

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

if -6s > -7s, what must be true about s?

OpenStudy (anonymous):

Solve for B: ax-by=c

OpenStudy (anonymous):

V=1/3Bh If h=7 and V =84, find B

OpenStudy (anonymous):

36

OpenStudy (anonymous):

thanks. could u break it down for me in dummy term :-). I'm slow over here tonight for some reason

OpenStudy (anonymous):

84 x 3 = B x 7 12 X 3 = B

OpenStudy (anonymous):

ok, thanks. now I get it

OpenStudy (anonymous):

one number is 3 less than 6 times the second number. Their sum is 46. Find the numbers.

OpenStudy (anonymous):

hi always start new post with new question, it will help others to see you question x + 6x -3 = 46 43/7 279/7

OpenStudy (anonymous):

Thanks for that advice. I appreciate all your help. I'm having a mental block over here :-)

OpenStudy (anonymous):

be cool dont worry you will be good with practice